# thanks: Corresponding author

### v1

@@ 4 Discussion
This article discusses over-luminous SNe Ia and their potential use in understanding and resolving Hubble tension. The use of over-luminous SNe Ia as standard candles affect the calibration of the distance ladder and hence can change the value of the Hubble constant. Using over-luminous SNe Ia leads to underestimating distances, it is expected to bring the value of the Hubble constant down from $H_{0}=73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ as measured in the SH0ES survey nearing to that of the early-Universe measurement. This study analyzes eight over-luminous SNe Ia listed in Table 1 and uses each of them as a standard candle to obtain corresponding values for the Hubble constant.
Utilizing a Bayesian statistical framework and a sample of eight over-luminous SNe Ia, we demonstrate a reduction in $H_{0}$ from $73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ to $68.686\pm 0.309\rm\,km\,s^{-1}\,Mpc^{-1}$ when using a flat prior. Alternatively, we obtain $69.373\pm 0.287\rm\,km\,s^{-1}\,Mpc^{-1}$ when employing a Gaussian prior based on Cepheid variable star measurements to estimate distances and the Hubble’s constant. These results exhibit closer concordance with CMB data, thereby prompting a critical reassessment of the viability of SNe Ia as reliable standard candles for cosmological distance determination. The observed luminosity of over-luminous SNe Ia suggests their potential utility in extending the observable cosmological horizon.
@@
However, this analysis introduces a novel discrepancy: the significant divergence between $H_{0}$ values obtained from over-luminous SNe Ia and those derived from regular SNe Ia. This raises fundamental questions regarding the intrinsic standardization of SNe Ia luminosities, particularly in light of the inclusion of over-luminous events. The application of over-luminous SNe Ia as distance indicators may lead to systematic underestimation of cosmological distances, contributing to their exclusion from standard candle applications. Furthermore, over-luminous SNe Ia exhibit substantial variability in peak luminosity, spectral characteristics, and light curve morphology. The potential existence of super-Chandrasekhar mass WD progenitors necessitates a rigorous reevaluation of standardization methodologies for over-luminous SNe Ia, analogous to those applied to standard events. A comprehensive theoretical understanding of the explosion mechanisms associated with super-Chandrasekhar mass WDs, as suggested in Das and Mukhopadhyay (2013a), may facilitate the development of over-luminous SNe Ia as a distinct class of standardizable candles. A primary limitation of this study is the restricted sample size, constrained by the lack of available data. Future investigations employing an expanded catalog of over-luminous SNe Ia have the potential to refine the $H_{0}$ estimate and contribute substantively to the resolution of the Hubble tension.
@@ Acknowledgments
TRG gratefully acknowledges Dr. Ravi Sheth from the University of Pennsylvania for insightful discussions and valuable input. AR and SK would like to thank The Institute of Mathematical Sciences (IMSc), Chennai for providing support and resources during the course of this study.

### v2

@@ 4 Discussion
This article discusses over-luminous SNe Ia and their potential use in understanding and resolving the Hubble tension. The use of over-luminous SNe Ia as standard candles affects the calibration of the distance ladder, and hence, can change the value of $H_{0}$. Using over-luminous SNe Ia leads to underestimating distances, and it is expected to bring the value of $H_{0}$ from $73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ as measured in the SH0ES survey, near that of the early-universe measurement. This study analyzes eight over-luminous SNe Ia listed in Table 1 to estimate $\mathsf{M_{B}}$ reported in previous studies with the same estimated using the Tripp formula. The goal is to understand the difference in the $\mathsf{M_{B}}$ values and determine $H_{0}$ for which the difference is minimized.
Utilizing a Bayesian statistical framework and a sample of eight over-luminous SNe Ia and eight normal luminosity SNe Ia, we demonstrate a reduction in $H_{0}$ from $73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ to $H_{0}=68.100\pm 1.048\rm\,km\,s^{-1}\,Mpc^{-1}$ with a flat prior and $H_{0}=71.518\pm 0.604\rm\,km\,s^{-1}\,Mpc^{-1}$ with the Gaussian prior based on Cepheid variable star measurements to estimate the distances. Moreover, combining them with standard luminosity SNe Ia, we achieve $H_{0}=70.622\pm 1.087\rm\,km\,s^{-1}\,Mpc^{-1}$ when using a flat prior. The Bayesian analysis yields significant deviation in $H_{0}$ estimate using over-luminous SNe Ia alone with repsct to the value reported by SH0ES collaboration, while a modest reduction in $H_{0}$ is obtained combining both over-luminous and standard one. This suggests that over-luminous SNe Ia can be reliably incorporated alongside standard SNe Ia in cosmological analyses for distance estimation. Nonetheless, the observed deviation underscores the need to extend the Phillips relation to account for over-luminous SNe Ia explicitly.
@@
These results point toward the possibility that modifications to the standard FLRW metric may be necessary to resolve the persistent tension in the measurement of $H_{0}$. A detailed discussion of such a modification is provided in Krishnan et al. (2021), which demonstrates a reduction in $H_{0}$ to $71\rm\,km\,s^{-1}\,Mpc^{-1}$, a value that closely aligns with our Bayesian result based solely on over-luminous SNe Ia. Table 2 shows an increase in the distance modulus with a decrease in the value of $H_{0}$. This leads to the conclusion that the distances measured using the standard luminosity SNe Ia are an underestimation, and corrections are required to obtain the correct distances. However, this analysis introduces a novel discrepancy: the significant divergence between $H_{0}$ values obtained from over-luminous SNe Ia and those derived from regular SNe Ia. This raises fundamental questions regarding the intrinsic standardization of SNe Ia luminosities, particularly considering the inclusion of over-luminous events. The application of over-luminous SNe Ia as distance indicators may lead to a systematic underestimation of cosmological distances, contributing to their exclusion from standard candle applications. Furthermore, over-luminous SNe Ia exhibit substantial variability in peak luminosity, spectral characteristics, and light curve morphology. The potential existence of super-Chandrasekhar mass WD progenitors necessitates a rigorous re-evaluation of standardization methodologies for over-luminous SNe Ia, analogous to those applied to standard events. A comprehensive theoretical understanding of the explosion mechanisms associated with super-Chandrasekhar mass WDs, as suggested in Das and Mukhopadhyay (2013a), may facilitate the development of over-luminous SNe Ia as a distinct class of standardizable candles. A few attempts have been made to explain the WD to SN Ia mechanism in the single degenerate (WD accreting matter from a second star) case in Chen and Li (2009); Hachisu et al. (2012). A primary limitation of this study is the small sample size, which was constrained by the lack of available data. Future investigations employing an expanded catalog of over-luminous SNe Ia have the potential to refine the $H_{0}$ estimate and contribute substantially to the resolution of Hubble tension.
@@ Acknowledgments
T.R.G. gratefully acknowledges Dr. Ravi Sheth from the University of Pennsylvania and Dr. S. Kalyana Rama from IMSc for insightful discussions and valuable comments. S.K. acknowledges funding by the National Center for Science, Poland, grant no. 2023/49/B/ST9/02777. A.R. and S.K. would like to thank IMSc, Chennai for providing support and resources during the course of this study.